Fermi National Laboratory
 

Operation Summary for June 15, 2011 - June 17, 2011

 

Wednesday June 15

The day shift began with Operations monitoring the Tevatron (TEV) store 8818 (345.59E30) and the Recycler stash (14.1E10), with the Antiproton Source (Pbar) stacking (38.1E10), with the Main Injector (MI), Linac, and Booster operating normally, with the NuMI, MiniBooNE, and Meson FTBF T-978 experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams.

During the day shift, at 10 AM, operators reported that the Linac quadrupole (Q7) bulk power supply tripped off; Linac personnel had the supply back on by 11 AM. MECAR memory issues required operators to reboot MECAR and let it relearn all cycles; this kept MI off for 30 minutes.

During the evening shift, TEV personnel worked on the Tevatron Electron Lens (TEL2) BPM front end from 5 PM to 7 PM. Operations disabled beam to T-978 at 6:04 PM. Operators reported some minor downtime due to a Booster RF station (BRF6).

Thursday June 16

The midnight shift began with Operations monitoring store 8818 (73.4E30) and the stash (361.4E10), with Pbar stacking (17.9E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, at 12:55 AM, MECAR caused a MI beam permit trip during a coalescing check; the Recycler lost 2E10 of its stash while MECAR was relearning the cycles. Operations terminated store 8818 at 1:48 AM, put the TEV into shot setup, and established store 8819 at 2:59 AM with an average initial luminosity (AIL) of 402.22E30. At 5:03 AM, operators reported more MECAR related MI permit trips that required MECAR to take the time to relearn. Operators reported some short downtimes due to BRF6, BRF7, LRF1, and due to a Linac Klystron RF station (KRF1).

During the day shift, operators turned off the MiniBooNE beam at 8 AM to perform a two-hour pressure test on the MI-12 service building; MiniBooNE resumed taking beam. Operators reported short downtimes due to LRF1, a LRF5 vacuum pump, MRF2, BRF1, and BRF6.

During the evening shift, at 4:23 PM, the MS2 water-to-air cooling tower temperature came into alarm. FESS personnel investigated and found the belts had come off of the motor for the section A cooling tower and the section B tower motor would not turn on. They replaced the belts for section A and determined that the motor for section B might need to be replaced; further investigation will be necessary. Operations disabled beam to T-978 at 5:59 PM. Operations terminated store 8819 at 6:28 PM, put the TEV into shot setup, and established store 8820 at 7:47 PM with an AIL of 399.18E30. TEV personnel turned on TEL2 for hollow beam collimation at 8:05 PM. Operators reported some short downtimes due to a Linac quadrupole power supply (QPS105), KRF4, and BRF7.

Friday June 17

The midnight shift began with Operations monitoring store 8820 (174.97E30) and the stash (140.0E10), with Pbar stacking (20.6E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, Operations established beam to T-978 at 4:30 AM. At 6:17 AM, operators reported that the Pbar Debuncher Lithium Lens tripped off on a LCW low flow indication; it reset. Operators contacted a Pbar person who believed the flow meter was bad and needed replacing. Operators reported short downtimes due to BRF6, BRF7, and to MI8 collimator problems.

The Plans for Friday and the Weekend

The plans for today are to turn off the TEV so that cryo system personnel can repair the TEV sector F4 wet engine flywheel. The plans for the weekend are to stack, stash, and store. The Run Coordinator will allow D0 an access to make repairs after store termination, during the cryo system repairs. The Coordinator will also allow Pbar to shut down to replace the Lens flow meter and will allow TEV personnel to conduct a study.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 06/17/2011    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory